# Gradient Descent with Large Step Sizes: Chaos and Fractal Convergence Region

### v2

@@ B.1 Topological entropy
The system $\sigma$ is called the full-shift on two symbols. Despite its simple definition, this system is unpredictable and chaotic (see, e.g., Devaney and Eckmann,, 1987). In particular, periodic points are dense in the state space, and, there exists a trajectory that is dense in the state space, i.e., there is a single trajectory that can come arbitrarily close to any point. In terms of predictability, consider two points $\boldsymbol{s}=(s_{0}s_{1}\cdots),\boldsymbol{s}^{\prime}=(s_{0}^{\prime}s_{1}^{\prime}\cdots)$ that have the same first $m$ elements, but differ starting from the $(m+1)$th element. By the definition of the distance, we have $d(\boldsymbol{s},\boldsymbol{s}^{\prime})\leq\sum_{j=0}^{\infty}1/2^{m+j}=1/2^{m+1}$. However, for all $N\geq m$, we have $d(\sigma^{N}(\boldsymbol{s}),\sigma^{N}(\boldsymbol{s}))>1/2$. Therefore, even if two initial points are arbitrarily close to each other, one can not make any prediction on how close their trajectories will remain in the long term. This unpredictability stems form the richness of “distinct” trajectories. In fact, one can show that $h(\sigma)=\log 2>0$ (see, e.g., Vries,, 2014). Note, the topological entropy of the gradient descent in matrix factorization is at least $\log 3$ (Theorem 1).
@@ B.2 Box-counting dimension and unpredictability
There have been numerous investigations discussing how a non-integer fractal dimension implies unpredictability in dynamical systems (see, e.g., Tél,, 1990; Aguirre et al.,, 2009). Here we provide a brief introduction to this topic.
@@
where $N(\varepsilon)$ is the number of boxes of side length $\varepsilon$ needed to cover the set $S$. For a dynamical system $F\colon X\to X$ where $X\subset\mathbb{R}^{D}$ is the state space, let $D_{B}$ be the box-counting dimension of a basin boundary and $D$ be the topology dimension of the state space. Consider a collection of trajectories and randomly perturb their initial points by a scale $\varepsilon$. Let $f(\varepsilon)$ denote the fraction of the trajectories that converge to a different point, i.e., whose initial point lies in a different basin of attraction after the perturbation. Thus, $f(\varepsilon)$ can be roughly viewed as the chance of making an error in predicting the converged point when the precision in specifying the initial point is $\varepsilon$. In general, the following scaling relation holds (Grebogi et al., 1983a, ):
@@ Appendix C Diagonality of the target matrix
We show that in matrix factorization (1), one may assume without lose of generality that the target matrix is diagonal. This simplification is a standard technique that has been widely adopted in the literature.
@@ Appendix D Quotient dynamics of gradient descent
We will show that the gradient descent dynamics are fully captured by their evolution across the fibers of the map $T$. In other words, different initializations in the same fiber produce qualitatively identical trajectories. This reflects an inherent symmetry of the system. The quotient system factors out this symmetry and describes the fiber-wise dynamics. The term quotient dynamical system is borrowed from the theory of equivariant dynamical systems (see, e.g., Golubitsky and Stewart,, 2003).
@@ D.1 Quotient dynamical system
We first introduce two properties of the map $T$: (i) the preimage of any measure zero set has measure zero and (ii) the fiber of $T$ is generically a smooth manifold.
@@
By Ponomarev, (1987), it suffices to show the map $T$ is a submersion almost everywhere, i.e., the Jacobian of $T$ has rank two almost everywhere. Notice that
@@
For almost all $(z,w)\in T(\mathbb{R}^{2d})=\left\{(z,w)\in\mathbb{R}^{2}\colon w\geq 2|z+\mu|\right\}$, $T^{-1}(z,w)$ is diffeomorphic to $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$.
@@
$T^{-1}(z,w)=\left\{(p,q)\colon\|p\|_{2}^{2}=w+2(z+\mu),\left\lVert q\right\rVert_{2}^{2}=w-2(z+\mu)\right\}.$
Thus, the fiber is diffeomorphic to $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$ whenever $w\pm 2(z+\mu)\neq 0$. Note this only fails at a measure-zero set, which completes the proof. ∎
@@
We have that $(z_{t+1},w_{t+1})=f(z_{t},w_{t})$ holds for all $t\geq 0$. In particular, $(u_{t},v_{t})$ converges to $\mathcal{M}$ if and only if $(z_{t},w_{t})$ converges to $T(\mathcal{M})$, and it converges to $(\mathbf{0},\mathbf{0})$ if and only if $(z_{t},w_{t})$ converges to (y,0)(-y,0).
@@ D.2 Properties of the quotient dynamics
We identifies three invariant sets of the quotient system $F$. A set $S\subset\Omega$ is said to be an invariant set under $F$ if $F(S)\subset S$.
@@
$\kappa(4|\mu|)=4|\mu|(|\mu|-1+\nu)^{2}=4|\mu|(|\mu|-a)^{2}.$
Since $0<|\mu|\leq a$ and $0\leq a<1$, $\kappa(4|\mu|)\leq 4|\mu|$. Therefore, the image set of $\kappa([0,4|\mu|])$ is contained $0,4|\mu|]$. This means that the set $\left\{Q=4|\mu|\right\}$ is invariant under $F$, which completes the proof. ∎
@@
Notice that the singular value of $F$ lies in the set
@@
We now show that $F$ maps $\partial A_{0}$ bijectively to $\partial\Omega$. Since $A_{0}\subset\mathrm{cl}(A_{0})$, we have that $F(A_{0})=\Omega^{o}\subset F(\mathrm{cl}(A_{0}))$. Since $\mathrm{cl}(A_{0})$ is compact, therefore, $F|_{\mathrm{cl}(A_{0})}\colon\mathrm{cl}(A_{0})\to\Omega$ is proper, and hence is closed (see, e.g., Theorem 4.95. Lee,, 2000). Therefore, $F(\mathrm{cl}(A_{0}))$ is a closed set that contains $\Omega^{o}$. Hence, $\mathrm{cl}(\Omega^{o})=\Omega\subset F(\mathrm{cl}(A_{0}))$. Since $F(A_{0})=\Omega^{o}$, we have $\partial\Omega\subset F(\partial A_{0})$, which means $F|_{\partial A_{0}}$ is onto $\partial\Omega$. We next show it is also injective. By Lemma 11, we know $F$ maps $\left\{z=\nu-1\right\}$ to $\left\{w=2(z+\mu)\right\}$ and maps $\left\{w=-2(z+\mu)\right\}$ to itself. When $z=\nu-1$, the $w$-update under $F$ is given by
@@
By differentiation computation, we claim that the last equation holds when $\nu\in[0,1]$. Therefore, we prove that the maximum $Q$ value on the parabola $\iota\colon 1-w-3z^{2}-2z\mu-2\nu+w\nu+\nu^{2}=0$ is at most $6-4\nu$. Notice that all the singular values of $F$ is given by $\partial\Omega\cup F(\iota)$. But what we have shown and Lemma 12, we have
@@ E.1.1 Preliminary results
As the Jacobian of $F$ has full rank almost everywhere, the above set is a measure-zero set (Ponomarev,, 1987). Hence, almost all initializations with $Q>8$ diverge.
@@
Consider gradient descent with step $\eta$ in problem (2). Any global minimizer with $\|u\|^{2}+\|v\|^{2}\geq 2/\eta$ is an unstable minimizer, i.e., it repels orbits in its neighborhood. Consequently, initializations that converge to such as minimizer form a measure-zero set. Moreover, when $|\mu|>1$, i.e., $\eta|y|>1$, all global minimizers are unstable.
@@
Note $\tilde{F}_{2}$ is a piecewise linear continuous map with slope equal to $\pm 3$. Hence, the topological entropy of $\tilde{F}_{2}$ is equal to $\log 3$ (De Melo and Van Strien,, 2012, Corollary of Theorem 7.2,). For a univariate map on a compact interval, positive topological entropy implies Devaney chaotic (Theorem 3.13, Elaydi,, 2007). Due to the conjugacy, we have that $\tilde{F}$ is also Devaney chaotic. Also, conjugacy preserves topological entropy (Theorem 1.7, Ch.8, Robinson,, 1998). Therefore, $h(\tilde{F})=\log 3$.
We now show the existence of periodic orbit with any period. According to the Li-Yorke Theorem (Li and Yorke,, 1975), a sufficient condition is that there exists a point $x$ such that $\tilde{F}_{2}^{3}(x)\leq x<\tilde{F}_{2}(x)<\tilde{F}_{2}^{2}(x)$. Consider x=5/7x=-5/7. We have that $\tilde{F}_{2}(x)=-1/7$, $\tilde{F}_{2}^{2}(x)=3/7$, and $\tilde{F}_{2}^{3}(x)=-5/7$. Due to the conjugacy, $\tilde{F}$ also has periodic orbit with any period and is Li-Yorke chaotic. This completes the proof. ∎
@@
Assume $|\mu|<1$. We have that $h(\mathrm{GD}_{\eta})\geq h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})\geq\log 3$, and $\mathrm{GD}_{\eta}$ admits periodic orbits with any period.
@@
By Lemma 12, $\partial\mathcal{D}_{\eta}^{\prime}$ is invariant under $\mathrm{GD}_{\eta}$. Since the map $T(u,v)=(z,w)$ is a semi-conjugacy between $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}}$ and $\tilde{F}$, we have that $h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})\geq h(\tilde{F})=\log 3$ (see, e.g., Theorem 1.7, Ch 8. Robinson,, 1998), where the equality comes from Proposition 17. Meanwhile, since $\partial\mathcal{D}_{\eta}^{\prime}$ is an invariant subset of $\mathbb{R}^{2d}$, we have that $h(\mathrm{GD}_{\eta})\geq h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}})$ (see, e.g., Proposition 8.1.7. Vries,, 2014).
@@
Note $\tilde{F}_{2}$ is a piecewise linear map defined on [1,1][-1,1], with each piece mapping onto the whole interval [1,1][-1,1]. With classical analyses in symbolic dynamics (see, e.g., Ch 1.7. Devaney and Eckmann,, 1987), we have that $\tilde{F}_{2}$, when restricted to $1,1]-\{x\in[-1,1]\colon\tilde{F}_{2}^{N}(x)\neq\pm 1,\forall N\}$, is conjugate to the shift operator $\sigma$, defined on the following set
@@
When $d\geq 2$, the system $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}}$ is not topological transitive.
@@
Let $(u_{k},v_{k})_{k\geq 0}$ be an orbit under $\mathrm{GD}_{\eta}$. As shown in (21), $u_{k}+v_{k}$ is always parallel to $u_{0}+v_{0}$. Consider the map $\tau\colon\mathbb{R}^{2d}\to\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$:
$\tau(u,v)=\left(\frac{u+v}{\|u+v\|},-\frac{u+v}{\|u+v\|}\right).$
For any open set $A\subset\mathbb{R}^{2d}$, we have $\tau(A)=\tau(\mathrm{GD}^{N}_{\eta}(A))$ for all $N\geq 0$. Since $d\geq 2$, we can always choose two open sets $A,B\subset\mathbb{R}^{2d}$ small enough such that $\tau(A)\cap\tau(B)=\varnothing$. This means that $\tau(\mathrm{GD}^{N}_{\eta}(A))\cap\tau(\mathrm{GD}^{N}_{\eta}(B))=\varnothing$ for all $N\geq 0$. Since $\tau(\mathrm{GD}^{N}_{\eta}(A)\cap\mathrm{GD}^{N}_{\eta}(B))\subset\tau(\mathrm{GD}^{N}_{\eta}(A))\cap\tau(\mathrm{GD}^{N}_{\eta}(B))$, we have the former set is also empty. Hence, $\mathrm{GD}^{N}_{\eta}(A)\cap\mathrm{GD}^{N}_{\eta}(B)$ is empty. Therefore, $\mathrm{GD}^{N}$ is not topological transitive. This gives the claimed result. ∎

### v3

@@ B.1 Topological entropy
The system $\sigma$ is called the full-shift on two symbols. Despite its simple definition, this system is unpredictable and chaotic (see, e.g., Devaney and Eckmann, 1987). In particular, periodic points are dense in the state space, and, there exists a trajectory that is dense in the state space, i.e., there is a single trajectory that can come arbitrarily close to any point. In terms of predictability, consider two points $\boldsymbol{s}=(s_{0}s_{1}\cdots),\boldsymbol{s}^{\prime}=(s_{0}^{\prime}s_{1}^{\prime}\cdots)$ that have the same first $m$ elements, but differ starting from the $(m+1)$th element. By the definition of the distance, we have $d(\boldsymbol{s},\boldsymbol{s}^{\prime})\leq\sum_{j=0}^{\infty}1/2^{m+j}=1/2^{m+1}$. However, for all $N\geq m$, we have $d(\sigma^{N}(\boldsymbol{s}),\sigma^{N}(\boldsymbol{s}))>1/2$. Therefore, even if two initial points are arbitrarily close to each other, one can not make any prediction on how close their trajectories will remain in the long term. This unpredictability stems from the richness of “distinct” trajectories. In fact, one can show that $h(\sigma)=\log 2>0$ (see, e.g., Vries, 2014). Note, the topological entropy of the gradient descent in matrix factorization is at least $\log 3$ (Theorem 1).
@@ B.2 Box-counting dimension and unpredictability
There have been numerous investigations discussing how a non-integer fractal dimension implies unpredictability in dynamical systems (see, e.g., Tél, 1990; Aguirre et al., 2009). Here we provide a brief introduction to this topic.
@@
where $N(\varepsilon)$ is the number of boxes of side length $\varepsilon$ needed to cover the set $S$. For a dynamical system $F\colon X\to X$ where $X\subset\mathbb{R}^{D}$ is the state space, let $D_{B}$ be the box-counting dimension of a basin boundary and $D$ be the topology dimension of the state space. Consider a collection of trajectories and randomly perturb their initial points by a scale $\varepsilon$. Let $f(\varepsilon)$ denote the fraction of the trajectories that converge to a different point, i.e., whose initial point lies in a different basin of attraction after the perturbation. Thus, $f(\varepsilon)$ can be roughly viewed as the chance of making an error in predicting the converged point when the precision in specifying the initial point is $\varepsilon$. In general, the following scaling relation holds (Grebogi et al., 1983a):
@@ Appendix C Diagonality of the target matrix
We show that in matrix factorization (1), one may assume without loss of generality that the target matrix is diagonal. This simplification is a standard technique that has been widely adopted in the literature.
@@ Appendix D Quotient dynamics of gradient descent
We will show that the gradient descent dynamics are fully captured by their evolution across the fibers of the map $T$. In other words, different initializations in the same fiber produce qualitatively identical trajectories. This reflects an inherent symmetry of the system. The quotient system factors out this symmetry and describes the fiber-wise dynamics. The term quotient dynamical system is borrowed from the theory of equivariant dynamical systems (see, e.g., Golubitsky and Stewart, 2003).
@@ D.1 Quotient dynamical system
We first introduce two properties of the map $T$: (i) the preimage of any measure-zero set has measure zero and (ii) the fiber of $T$ is generically a smooth manifold.
@@
By Ponomarev (1987), it suffices to show the map $T$ is a submersion almost everywhere, i.e., the Jacobian of $T$ has rank two almost everywhere. Notice that
@@
For almost all $(z,w)\in T(\mathbb{R}^{2d})=\left\{(z,w)\in\mathbb{R}^{2}\colon w\geq 2|z+y|\right\}$, $T^{-1}(z,w)$ is diffeomorphic to $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$.
@@
$T^{-1}(z,w)=\left\{(p,q)\colon\|p\|_{2}^{2}=w+2(z+y),\left\lVert q\right\rVert_{2}^{2}=w-2(z+y)\right\}.$
Thus, the fiber is diffeomorphic to $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}$ whenever $w\pm 2(z+y)\neq 0$. Note this only fails at a measure-zero set, which completes the proof. ∎
@@
We have that $(z_{t+1},w_{t+1})=f(z_{t},w_{t})$ holds for all $t\geq 0$. In particular, $(u_{t},v_{t})$ converges to $\mathcal{M}$ if and only if $(z_{t},w_{t})$ converges to $T(\mathcal{M})$, and it converges to $(\boldsymbol{0},\boldsymbol{0})$ if and only if $(z_{t},w_{t})$ converges to (y,0)(-y,0).
@@ D.2 Properties of the quotient dynamics
We identify three invariant sets of the quotient system $F$. A set $S\subset\Omega$ is said to be an invariant set under $F$ if $F(S)\subset S$.
@@
$\kappa(4|\mu|)=4|\mu|(|\mu|-1+\nu)^{2}.$
Since $\nu<1-|\mu|$, we have that $1<|\mu|-1+\nu<0$ and that $\kappa(4|\mu|)\leq 4|\mu|$. Therefore, the image set of $\kappa([0,4|\mu|])$ is contained $0,4|\mu|]$. This means that the set $\left\{Q=4|\mu|\right\}$ is invariant under $F$, which completes the proof. ∎
@@
Notice that the critical points of $F$ lie in the set
@@
We now show that $F$ maps $\partial A_{0}$ bijectively to $\partial\Omega$. Since $A_{0}\subset\mathrm{cl}(A_{0})$, we have that $F(A_{0})=\Omega^{o}\subset F(\mathrm{cl}(A_{0}))$. Since $\mathrm{cl}(A_{0})$ is compact, therefore, $F|_{\mathrm{cl}(A_{0})}\colon\mathrm{cl}(A_{0})\to\Omega$ is proper, and hence is closed (see, e.g., Lee, 2000, Theorem 4.95). Therefore, $F(\mathrm{cl}(A_{0}))$ is a closed set that contains $\Omega^{o}$. Hence, $\mathrm{cl}(\Omega^{o})=\Omega\subset F(\mathrm{cl}(A_{0}))$. Since $F(A_{0})=\Omega^{o}$, we have $\partial\Omega\subset F(\partial A_{0})$, which means $F|_{\partial A_{0}}$ is onto $\partial\Omega$. We next show it is also injective. By Lemma 12, we know $F$ maps $\left\{z=\nu-1\right\}$ to $\left\{w=2(z+\mu)\right\}$ and maps $\left\{w=-2(z+\mu)\right\}$ to itself. When $z=\nu-1$, the $w$-update under $F$ is given by
@@
By differentiation computation, we claim that the last equation holds when $\nu\in[0,1]$. Therefore, we prove that the maximum $Q$ value on the parabola $\iota\colon 1-w-3z^{2}-2z\mu-2\nu+w\nu+\nu^{2}=0$ is at most $6-4\nu$. Notice that all the critical values of $F$ is given by $\partial\Omega\cup F(\iota)$. But what we have shown and Lemma 13, we have
@@ E.1.1 Preliminary results
As the Jacobian of $F$ has full rank almost everywhere, the above set is a measure-zero set (Ponomarev, 1987). Hence, almost all initializations with $Q>8$ diverge.
@@
Consider gradient descent with step $\eta$ in problem (2). Any global minimizer with $\|u\|^{2}+\|v\|^{2}\geq 2/\eta$ is an unstable minimizer, i.e., it repels orbits in its neighborhood. Consequently, initializations that converge to such a minimizer form a measure-zero set. Moreover, when $|\mu|>1$, i.e., $\eta|y|>1$, all global minimizers are unstable.
@@
Note $\tilde{F}_{2}$ is a piecewise linear continuous map with slope equal to $\pm 3$. Hence, the topological entropy of $\tilde{F}_{2}$ is equal to $\log 3$ (De Melo and Van Strien, 2012, Corollary of Theorem 7.2). For a univariate map on a compact interval, positive topological entropy implies Devaney chaotic (Elaydi, 2007, Theorem 3.13). Due to the conjugacy, we have that $\tilde{F}$ is also Devaney chaotic. Also, conjugacy preserves topological entropy (Robinson, 1998, Theorem 1.7, Ch.8). Therefore, $h(\tilde{F})=\log 3$.
We now show the existence of periodic orbit with any period. According to the Li-Yorke Theorem (Li and Yorke, 1975), a sufficient condition is that there exists a point $x$ such that $\tilde{F}_{2}^{3}(x)\leq x<\tilde{F}_{2}(x)<\tilde{F}_{2}^{2}(x)$. Consider x=5/7x=-5/7. We have that $\tilde{F}_{2}(x)=-1/7$, $\tilde{F}_{2}^{2}(x)=3/7$, and $\tilde{F}_{2}^{3}(x)=-5/7$. Due to the conjugacy, $\tilde{F}$ also has periodic orbit with any period and is Li-Yorke chaotic. This completes the proof. ∎
@@
Assume $|\mu|<1$. We have that $h(\mathrm{GD}_{\eta})\geq h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}^{\prime}})\geq\log 3$, and $\mathrm{GD}_{\eta}$ admits periodic orbits with any period.
@@
By Lemma 13, $\partial\mathcal{D}_{\eta}^{\prime}$ is invariant under $\mathrm{GD}_{\eta}$. Since the map $T(u,v)=(z,w)$ is a semi-conjugacy between $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}^{\prime}}$ and $\tilde{F}$, we have that $h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}^{\prime}})\geq h(\tilde{F})=\log 3$ (see, e.g., Robinson, 1998, Theorem 1.7, Ch 8), where the equality comes from Proposition 18. Meanwhile, since $\partial\mathcal{D}_{\eta}^{\prime}$ is an invariant subset of $\mathbb{R}^{2d}$, we have that $h(\mathrm{GD}_{\eta})\geq h(\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}^{\prime}})$ (see, e.g., Vries, 2014, Proposition 8.1.7).
@@
Note $\tilde{F}_{2}$ is a piecewise linear map defined on [1,1][-1,1], with each piece mapping onto the whole interval [1,1][-1,1]. With classical analyses in symbolic dynamics (see, e.g., Devaney and Eckmann, 1987, Ch 1.7), we have that $\tilde{F}_{2}$, when restricted to $1,1]-\{x\in[-1,1]\colon\tilde{F}_{2}^{N}(x)\neq\pm 1,\forall N\}$, is conjugate to the shift operator $\sigma$, defined on the following set
@@
Assume that $\eta|y|<1$ and $d\geq 2$. The system $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}^{\prime}}$ is not topological transitive.
@@
Let $(u_{k},v_{k})_{k\geq 0}$ be an orbit under $\mathrm{GD}_{\eta}$. Consider new coordinates $p_{k}=u_{k}+v_{k}$ and $q_{k}=u_{k}-v_{k}$ for $k\geq 0$. Recall (22) and notice that $z_{k}=u_{k}^{\top}v_{k}-y=-y+(\|p_{k}\|^{2}-\|q_{k}\|^{2})/4$. Thus, $(p_{k},q_{k})$ evolves autonomously. Since $(u,v)\mapsto(p,q)$ is a homeomorphism, the $uv$-system is topological transitive if and only if the $pq$-system is.
Notice that, if $4=\eta(\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2})-\eta^{2}y(u^{\top}v-y)$, we have that
$4\geq\eta(\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2})-\eta^{2}|y|\frac{\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2}}{2}=\eta(1-\frac{\eta|y|}{2})(\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2}),$
where we used that $|u^{\top}v|\leq(\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2})/2$. Since $\eta|y|<1$, we have that $\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2}<8/\eta$. Therefore,
$(u,v)\in\partial\mathcal{D}_{\eta}^{\prime}$ $\Leftrightarrow\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2}+\sqrt{(\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2})^{2}-16y(u^{\top}v-y)}=\frac{8}{\eta}$ (23)
$\Leftrightarrow 4=\eta(\left\lVert u\right\rVert_{2}^{2}+\left\lVert v\right\rVert_{2}^{2})-\eta^{2}y(u^{\top}v-y)$
$\Leftrightarrow(u^{\top}\ v^{\top})\begin{pmatrix}\eta I_{d}&-\eta^{2}y/2\cdot I_{d}\\ -\eta^{2}y/2\cdot I_{d}&\eta I_{d}\end{pmatrix}\begin{pmatrix}u\\ v\end{pmatrix}=4-\eta^{2}y^{2}.$
The last equation is a quadratic form and the eigenvalues of the coefficient matrix are $\eta\pm\frac{1}{2}\eta^{2}y$, each with multiplicity $d$. Therefore, when $\eta|y|<1$, the quadratic form is positive definite and defines a smooth ellipsoid of dimension 2d12d-1. Notice that $(u,v)\mapsto(p,q)=(u+v,u-v)$ is a bijective linear transformation. Thus, $\partial\mathcal{D}_{\eta}^{\prime}$ is still an ellipsoid in $pq$-coordinates.
Now fix any $(p_{0},q_{0})\in\partial\mathcal{D}_{\eta}^{\prime}$ and an open neighborhood $U=U_{p}\times U_{q}$ of $(p_{0},q_{0})$ where $p_{0}\in U_{p}\subset\mathbb{R}^{d}$ and $q_{0}\in U_{q}\subset\mathbb{R}^{d}$. According to (22), we have that
$\cup_{k=0}^{\infty}\mathrm{GD}_{\eta}^{k}(U)\subset\mathcal{C}\triangleq\left\{cp\colon p\in U_{p},c\in\mathbb{R}\right\}\times\left\{dq\colon q\in U_{q},d\in\mathbb{R}\right\}\subset\mathbb{R}^{2d}.$
Clearly, when $d\geq 2$ and $U$ is small enough such that $(\boldsymbol{0},\boldsymbol{0})\notin U$, $\mathcal{C}$ is not dense in $\mathbb{R}^{2d}$ and there exists an open set $V$ such that $V\cap\partial\mathcal{D}_{\eta}^{\prime}\neq\varnothing$ and $\mathcal{C}\cap V=\varnothing$. Therefore, $\mathrm{GD}_{\eta}|_{\partial\mathcal{D}_{\eta}^{\prime}}$ is not transitive in $pq$-coordinates. This completes the proof. ∎